OpenStudy (anonymous):

Solve –4(x + 3) = 16

jimthompson5910 (jim_thompson5910):

-4(x+3)=16 -4x-12=16 -4x=16+12 -4x=28 x=28/(-4) x=-7

OpenStudy (anonymous):

\[-4(x+3)=16\] \[-4(x) +(-4)(3) = 16\] \[-4x-12=16\] \[-4x=16+12\] \[-4x=28\] \[x=28/(-4)\] \[x=-7\]